OpenStudy (anonymous):

Simplify... (5a^3 + 6a) + (-3a^2 - 5a)

OpenStudy (anonymous):

step 1 ... take off bracket.

OpenStudy (anonymous):

and you'll have: \(\LARGE 5a^3+6a-3a^2-5a\)

OpenStudy (anonymous):

now what can we do ? :)

OpenStudy (anonymous):

@Kreshnik 5a^3 + 3a^2 - a

OpenStudy (anonymous):

well, quite close. Just sign of \(\LARGE 3a^2\) must be negative and sign of "a" is positibe

OpenStudy (anonymous):

because you have: \[\LARGE 5a^3+6a-3a^2-5a=\] \[\LARGE 5a^3+(6-5)a-3a^2=\] \[\LARGE 5a^3+1a-3a^2=\] \[\LARGE 5a^3+a-3a^2=\]

OpenStudy (anonymous):

@Kreshnik Thanks

OpenStudy (anonymous):

no sign of 3a2 must be negative becoz (-3a)^2 will be 3a2

OpenStudy (anonymous):

@cesar2462 My pleasure xD @d.t. note.. you have \(\LARGE -3a^2\) not \(\LARGE (-3a)^2\)

OpenStudy (anonymous):

according to you... @d.t. we wouldn't even have 3a^2 it would be 9a^2 ! :)

OpenStudy (anonymous):

oh ...i got it thanks

OpenStudy (anonymous):

Nvm

OpenStudy (anonymous):

Anyone